OpenStudy (anonymous):

A raffle has a grand prize of a Caribbean cruise valued at $6000 with a second prize of a resort vacation valued at $1400. If each ticket costs $5 and 10400 tickets are sold, what are the expected winnings for a ticket buyer? Express to at least three decimal place accuracy in dollar form (as opposed to cents).

OpenStudy (anonymous):

Your chance of winning first prize = 1/10400. Your chance of winning second prize = 1/10399. Your chance of winning 1st or second prize is the sum of those two. Expected winning = 6000/10400 + 1400/10399.

OpenStudy (anonymous):

it is incorrect though!!

OpenStudy (anonymous):

You need to subtract $5 ticket price from it.

OpenStudy (anonymous):

Thanks.I got it!!also,the prob. of wining nothing!!